An amount of Rs 735 are divided among A,B and C.If each of them receives Rs 25 less,their shares would have been in the ratio 1:3:2.The money received by C was?
Options are
A. Rs 195
B. Rs 200
C. Rs 225
D. Rs 245

Answer by edjones(8007)

735-25*3=660
660/(1+3+2)=110
110*2=220
220+25=Rs 245 money received by C.
.
Ed